can you PLEASE help me!

Solve the systems of linear equations using the substitution method:
x=y/2
4y=3x-5

Found 2 solutions by Alan3354, Mathtut:
Answer by Alan3354(69443)

can you PLEASE help me!
Solve the systems of linear equations using the substitution method:
x=y/2
4y=3x-5
---------
No reason to panic.
Since eqn 1 already has x in terms of y, sub for x in eqn 2.
4y=3x-5
4y = 3(y/2) - 5
8y = 3y - 10
5y = -10
y = -2
----------
x = y/2
x = -1

Answer by Mathtut(3670)

x=y/2...........eq 1
4y=3x-5.........eq 2
:
take the value of x in equation 1 and plug it into every occurence of x in eq 2
:
4y=3(y/2)-5
:
8y=3y-10.....multiplied all terms by 2 to eliminate fraction
:
5y=-10
:
y=-2
:
now take y's found value and plug it back into either equation to find x. I choose eq 2
:
4(-2)=3x-5--->3x=-3
:
x=-1